A rectangle with a length more than four times the width has an area of 132km^2. Find x.
Answer by tommyt3rd(5050)

Without exact values you certainly cannot expect an exact answer...however if length is exactly 4 times width...
l=4x
A=lx=4x^2=132
x=sqrt(132/4)=sqrt(33)